Given that Kw for water is 2.4× 10–14 at 37 °C, compute the pH of a neutral aqueous solution at 37 °C, which is the normal human body temperature.
Answer #1

Kw for water = 2.4× 10^-14

Kw = [H+] x [OH-]

for ionization of water . [H+] = [OH-].

then Kw = [H+]^2

[H+]^2 = 2.4× 10^-14

[H+] = 1.55 x 10^-7 M

Now :

pH = -log [H+] = - log (1.55 x 10^-7 )

pH = 6.8